OpenStudy (anonymous):

The 18th term of a sequence is 243, and the 90th term is 516.6. What is the 125th term?

OpenStudy (anonymous):

do we know if this is a geometric or arithmetic sequence?

OpenStudy (anonymous):

guess it depends on what kind of sequence it is

OpenStudy (saifoo.khan):

Right.

OpenStudy (anonymous):

or neither o.O

OpenStudy (anonymous):

I'm not exactly sure but the options are: 624.6 649.6 653.4 4,928.4

OpenStudy (gw2011):

I get the 125th term as 649.6 This is the way I arrived at this answer: 90-18=72 terms between these two terms 516.6-243=273.6 is the increase between the 18th and 90th terms 273.6/72=3.8 which is the increase per term 125-18=107 terms between the 18th term and the 125th term (107)(3.8)=406.6 406.6+243=649.6